The largest swing in the world has a string length of 300 meters. If you are at a position at 60° with the vertical. What is your final velocity when you reach the bottom of the swing?
1
Expert's answer
2019-12-12T06:19:21-0500

From the conservation of energy:


"0.5mv^2=mgh"

"v^2=2gl(1-\cos{60})"

"v^2=2(9.8)(300)(1-0.5)"

"v=54\frac{m}{s}"